# Towards Enterprise-Ready Computer Using Generalist Agent

### v2

@@ 4 Results
Figure 5 illustrates the overall performance improvement achieved through the iterative evolution of our architecture. Each iteration involved two evaluations: first, a validation on the previously used sample to confirm the impact of enhancements aimed at addressing observed failures, and second, a test on a larger, more representative sample to assess generalizability. The description of the samples we used is detailed in Table 1. It is important to note that, in some instances, increasing the sample size resulted in a slight reduction in performance. This phenomenon reflects the inherent approximation of our sampling strategy, where smaller samples may not fully capture the complexity of the benchmark. Despite these minor fluctuations, the graph demonstrates a clear and consistent trend of improvement in task completion rate over time. This upward trajectory highlights the effectiveness of our iterative refinement process, emphasizing the value of continuous analysis and targeted enhancements in developing a robust and high-performing agentic system.
Table 1: Sample Sizes Used for Iterative Development
@@
Figure 5: Evolution of CUGA Architecture Performance. The graph illustrates the task completion rate (%) across different evaluation runs. Each run is denoted by the iteration and the dataset sample name. The size of the markers corresponds to the sample size where in each iteration we validate performance gains on the previous sample, as well as enlarge the sample to learn about new failures. Note that in some instances, increasing the sample size led to a minor reduction in performance, reflecting the approximation inherent in our sampling strategy. Overall, the graph demonstrates a clear trend of improvement in task completion rate over time, showcasing the effectiveness of our iterative refinement process.
@@ 5 Lessons Learned, Challenges, and Roadmap
Our iterative development methodology, coupled with smart evaluation and analytics tools, has been crucial for rapid progress and achieving state-of-the-art results cost-effectively. These tools have enabled us to effectively assess failures and identify areas for improvement. However, the process of finding, classifying, and aggregating root causes of failures within agentic trajectories remains labor-intensive, even with current tooling. Automating some of this manual effort, perhaps by leveraging AI agents, could significantly accelerate the development of agentic systems and potentially pave the way for fully autonomous evolution. Another interesting research opportunity lies in the area of smart regressions and how to evaluate individual agents within a full agentic system. Currently, best practices and lessons learned in this area are not widely shared. We hope this paper contributes to the community by sharing our experiences. Within IBM, we are leveraging insights, methodologies, and tools from initiatives like IBM CUGA and others, integrating and hardening them into the observability and analytics layers of the IBM Agentic Middleware platform. This will allow clients to benefit from these capabilities at a product-grade level.
@@
Creating and maintaining effective benchmarks is a significant challenge. Many existing benchmarks are developed within academic settings, reflecting the tasks and tools readily available to researchers. While valuable, these benchmarks often lack the complexity and nuances of real-world scenarios. A promising recent effort, TheAgentCompany [5], aims to address this by simulating more realistic tasks and incorporating human-agent communication. However, current benchmarks often focus on "happy path" scenarios and neglect critical aspects like safety, especially for agents interacting with enterprise systems, applications, and APIs. They typically do not evaluate agent behavior in exceptional circumstances or when human-in-the-loop interaction is required. A notable effort in this direction is STWebAgentBench [3], which extends WebArena by incorporating safety and policy adherence measurements alongside task completion. Ultimately, realistic benchmarks should mirror the specific environments, policies, and enterprise use cases of individual organizations. At IBM Research, we plan to further evolve CUGA to be a safe and policy-compliant, and to be a collaborative and trustworthy AI agent with human-in-the-loop support. Furthermore, we plan to evaluate CUGA on several leading benchmarks. These benchmarks combine multiple tools, desktop applications, coding, and APIs. We will also contribute to the community’s efforts in developing more comprehensive evaluation frameworks.
@@
Traditional system development relies on programming languages and low-code/no-code tools to build and test systems according to specifications defined in requirements documents. Agentic systems offer a radically different paradigm. Agents can potentially learn to perform tasks much like humans: by studying documentation and policies, receiving instructions, observing expert behavior, analyzing videos, and through interactive discovery and safe exploration. Emerging research suggests this vision is within reach. At IBM Research, we have begun experimenting with methods to enable CUGA to learn from unstructured documents and empower non-technical users to configure and customize its behavior to meet their specific needs.
@@
While our current state-of-the-art results are based on a frontier model, this choice was primarily driven by time efficiency, allowing us to rapidly iterate on our agentic architecture for the WebArena benchmark. We recognize the significant advantages of smaller and open models, particularly in terms of accessibility, affordability, privacy, efficiency, and cost. We have initiated experiments evaluating our architecture with these models and plan to publish a comprehensive evaluation of our findings upon completion.
@@ 6 Conclusion
In this paper, we have presented the current status of our work on developing IBM CUGA. We detailed the iterative evolution of our agentic architecture, the methodology and tools that facilitated rapid learning and cost-effective improvements, and key lessons learned and challenges ahead. Our results demonstrate the effectiveness of our approach, achieving state-of-the-art performance on the WebArena benchmark. We believe that our contributions provide valuable insights and practical guidance for the community, paving the way for future advancements in the field of enterprise-ready agentic AI systems.

### v3

@@ 4 Results
Figure 7 illustrates the overall performance improvement achieved through the iterative evolution of our architecture. Each iteration involved two evaluations: first, a validation on the previously used sample to confirm the impact of enhancements aimed at addressing observed failures, and second, a test on a larger, more representative sample to assess generalizability. The description of the samples we used is detailed in Table 2. It is worth noting that, in some instances, increasing the sample size has resulted in a slight reduction in performance. This phenomenon reflects the inherent approximation of our sampling strategy, where smaller samples may not fully capture the complexity of the benchmark. Despite these minor fluctuations, the graph demonstrates a clear and consistent trend of improvement in task completion rate over time. This upward trajectory highlights the effectiveness of our iterative refinement process, emphasizing the value of continuous analysis and targeted enhancements in developing a robust and high-performing agentic system.
Table 2: Progressively larger WebArena samples used in iterative development, enabling failure-driven refinement and generalization assessment across increasing benchmark coverage.
@@
Figure 7: Evolution of CUGA Architecture Performance. The graph illustrates the task completion rate (%) across different evaluation runs. Each run is denoted by the iteration and the dataset sample name. The size of the markers corresponds to the sample size, where in each iteration we validate performance gains on the previous sample, as well as enlarge the sample to learn about new failures. Note that in some instances, increasing the sample size led to a minor reduction in performance, reflecting the approximation inherent in our sampling strategy. Overall, the graph demonstrates a clear trend of improvement in task completion rate over time, showcasing the effectiveness of our iterative refinement process.
The AppWorld benchmark provides a structured and multi-layered evaluation of agents’ ability to interact with API-driven applications across increasing levels of complexity. As shown in Table 3, CUGA achieves a task goal completion rate of 73.2% under normal conditions and 57.6% in the more challenging settings. At the scenario level — which requires successful completion of all task instances within the same task template, usually 3 different instances — the agent reaches 62.5% and 48.2% respectively, reflecting strong generalization across diverse real-world tasks. Notably, CUGA is now the state-of-the-art performer on the AppWorld benchmark, ranking first on the official leaderboard.
Performance across levels reveals important trends. On Level 1 scenarios, which typically involve single app tasks, 2-3 API calls or linear workflows, the agent performs nearly flawlessly, achieving over 91% task goal completion and 84–87% scenario completion. This indicates the robustness of CUGA’s sub-agent decomposition and basic API reasoning capabilities. As complexity increases in Levels 2 and 3 - where tasks require variable tracking, conditional logic, and interaction across multiple APIs—performance degrades more gradually. The agent still maintains above 68% scenario completion on Level 2 (normal), but dips to 38.1–38.5% at Level 3, revealing key areas for further improvement in reasoning over branching workflows and handling unexpected outputs.
The average number of interactions per scenario also increases with difficulty, suggesting that CUGA adapts its strategy to the problem space, rather than simply following static plans. This is particularly evident in the challenge split, where the agent demonstrates a more efficient interaction pattern in early levels (fewer steps), but needs longer trajectories to resolve ambiguities or recover from partial failures in more complex scenarios. Overall, these results demonstrate that CUGA not only achieves strong performance in controlled web environments (WebArena), but also generalizes effectively to dynamic API-based tasks (AppWorld).
Table 3: CUGA performance on the AppWorld benchmark across increasing levels of API complexity, showing strong task and scenario completion rates and adaptive interaction strategies under both normal and challenge settings.
Level Task Goal Completion (%) Scenario Goal Completion (%) Avg. Interactions
Normal Challenge Normal Challenge Normal Challenge
All 73.2 57.6 62.5 48.2 10.69 8.40
Level 1 91.2 91.7 84.2 87.5 5.94 4.65
Level 2 77.1 58.7 68.8 42.0 10.36 8.33
Level 3 54.0 44.1 38.1 38.5 12.69 11.86
@@ 5 Lessons Learned, Challenges, and Roadmap
Our iterative development methodology, coupled with smart evaluation and analytics tools, has been crucial for rapid progress and achieving state-of-the-art results cost-effectively. These tools have enabled us to assess failures and identify areas for improvement effectively. However, the process of finding, classifying, and aggregating root causes of failures within agentic trajectories remains labor-intensive, even with current tooling. Automating some of this manual effort by leveraging AI agents could significantly accelerate the development of agentic systems and potentially pave the way for fully autonomous evolution. Another interesting research opportunity lies in the area of smart regressions and evaluating individual agents within a comprehensive agentic system. Currently, best practices and lessons learned in this area are not widely shared. We hope this paper contributes to the community by sharing our experiences. Within IBM, we are leveraging insights, methodologies, and tools from initiatives like IBM CUGA and others, integrating and hardening them into the observability and analytics layers of the IBM Agentic Middleware platform. This will allow clients to benefit from these capabilities at a product-grade level.
@@
Creating and maintaining effective benchmarks is a significant challenge. Many existing benchmarks are developed within academic settings, reflecting the tasks and tools readily available to researchers. While valuable, these benchmarks often lack the complexity and nuances of real-world scenarios. A promising recent effort, TheAgentCompany [6], aims to address this by simulating more realistic tasks and incorporating human-agent communication. However, current benchmarks often focus on "happy path" scenarios and overlook critical aspects, such as safety, particularly for agents interacting with enterprise systems, applications, and APIs. They typically do not evaluate agent behavior in exceptional circumstances or when human-in-the-loop interaction is required. A notable effort in this direction is STWebAgentBench [3], which extends WebArena by incorporating safety and policy adherence measurements alongside task completion. Ultimately, realistic benchmarks should mirror the specific environments, policies, and enterprise use cases of individual organizations.
At IBM Research, we plan to further evolve CUGA to be a safe and policy-compliant, as well as a collaborative and trustworthy AI agent with human-in-the-loop support. Furthermore, we plan to evaluate CUGA on several leading benchmarks. These benchmarks combine multiple tools, desktop applications, coding, and APIs. We will also contribute to the community’s efforts in developing more comprehensive evaluation frameworks. In particular, given the structure and scope of the AppWorld benchmark, we plan to explore the dimension of API safety, including how agents can reason about, restrict, and validate API usage in alignment with enterprise policies and security constraints. The AppWorld benchmark stands out as a remarkably well-designed and comprehensive evaluation suite, thoughtfully constructed to cover a wide range of realistic use cases across multiple applications. Its inclusion of diverse task types and multi-application workflows makes it a valuable resource for advancing the development of robust and enterprise-ready agentic systems.
@@
Traditional system development relies on programming languages and low-code/no-code tools to build and test systems according to specifications defined in requirements documents. Agentic systems offer a radically different paradigm. Agents can learn to perform tasks much like humans: by studying documentation and policies, receiving instructions, observing expert behavior, analyzing videos, and through interactive discovery and safe exploration. Emerging research suggests this vision is within reach. At IBM Research, we have begun experimenting with methods to enable CUGA to learn from unstructured documents, empowering non-technical users to configure and customize its behavior to meet their specific needs.
@@
While our current state-of-the-art results are based on a frontier model, this choice was primarily driven by time efficiency, allowing us to iterate on our agentic architecture across various benchmarks rapidly. We recognize the significant advantages of smaller and open models, particularly in terms of accessibility, affordability, privacy, efficiency, and cost. We have initiated experiments to evaluate our architecture using these models and plan to publish a comprehensive evaluation of our findings upon completion.
@@ 6 Conclusion
In this paper, we have presented the current status of our work on developing IBM CUGA. We detailed the iterative evolution of our agentic architecture, the methodology and tools that facilitated rapid learning and cost-effective improvements, and key lessons learned and challenges ahead. Our results demonstrate the effectiveness of our approach, achieving state-of-the-art performance on both the WebArena and AppWorld benchmarks. Our contributions offer valuable insights and practical guidance to the community, paving the way for future advancements in the field of enterprise-ready, agentic AI systems.